OpenStudy (anonymous):

Rationalize the denominator of 2/ ^3√a-1

OpenStudy (jdoe0001):

do you know what a "conjugate" is?

OpenStudy (solomonzelman):

Multiply top and bottom times, \(\large\color{blue}{ \rm \sqrt[3]{a-1} }\)

OpenStudy (anonymous):

Sort of

OpenStudy (jdoe0001):

\(\bf \Large \cfrac{2}{\sqrt[3]{a-1}}\quad ?\)

OpenStudy (anonymous):

Is it basically the opposite?

OpenStudy (anonymous):

Yes, that is the problem

OpenStudy (jdoe0001):

so... haemm ok.... well what would you get if you multiply top and bottom by \(\bf \sqrt[3]{(a-1)^2}\) ?

OpenStudy (anonymous):

1\[2\sqrt[3]{(a-1)^{2}}/a-1\]

OpenStudy (anonymous):

I think

OpenStudy (jdoe0001):

well one sec

OpenStudy (jdoe0001):

\(\bf \large{ \cfrac{2}{\sqrt[3]{a-1}}\cdot \cfrac{\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^2}}\implies \cfrac{2\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^1}\cdot \sqrt[3]{(a-1)^2}} \\ \quad \\ \cfrac{2\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^{1+3}}}\implies \cfrac{2\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^3}} \\ \quad \\ \cfrac{2\sqrt[3]{(a-1)^2}}{a-1} }\)

OpenStudy (jdoe0001):

hmm \(\bf \large { \cfrac{2}{\sqrt[3]{a-1}}\cdot \cfrac{\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^2}}\implies \cfrac{2\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^1}\cdot \sqrt[3]{(a-1)^2}} \\ \quad \\ \cfrac{2\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^{1+2}}}\implies \cfrac{2\sqrt[3]{(a-1)^2}}{\sqrt[3]{(a-1)^3}} \\ \quad \\ \cfrac{2\sqrt[3]{(a-1)^2}}{a-1} }\)

OpenStudy (jdoe0001):

anyhow just a typo =)

OpenStudy (jdoe0001):

the idea being, you'd multiply by a factor of the denominator, enough to get the radicand out

OpenStudy (anonymous):

Thanks! I was a little confused on the radicand part on the bottom and if it just canceled out or what.